# Axion Superradiance in Dipole Magnetic Fields of Pulsars

### v1

@@ I Introduction
The paper is organized as follows: In Sec. II we briefly introduce the calculational framework we use and explicit definitions and formulas needed for the evaluation of the final results. In Sec. III we present our main results and apply them to constrain axion parameter space using observations of MSPs. In Sec. IV we summarize and discuss further prospects of this work. Detailed intermediate results are provided in Appendices.
@@ II.2 Matrix elements for a dipole magnetic field
In [43] superradiance was demonstrated in the simple case of a magnetic field in the $\hat{\bf{z}}$-direction. This turns out to approximately capture the essential features of the system in the $l=|m|=1$ case, as the axion cloud is localized close to the equator, where the physical magnetic field is approximately in the $\hat{\bf{z}}$-direction. However, to extract more precise phenomenological predictions one must model the magnetic field more carefully. We therefore assume the more realistic dipole magnetic field
$\mathbf{B}(r,\theta,\phi)=\frac{B_{0}}{r^{3}}(2\cos\theta\hat{\bm{r}}+\sin\theta\hat{\bm{\theta}}),$ (21)
@@
$\left\langle\phi_{lmn}\left|V_{a\gamma\gamma}\right|A_{l^{\prime}m^{\prime}}^{(i)}\right\rangle$ $=(-1)^{-m}g_{\rm{a}\gamma\gamma}B_{0}R^{3}\mathcal{N}_{\phi}(\mu\alpha\omega)^{\frac{3}{2}}\mathcal{I}^{(i)}_{\Omega}I[l-2,l,l^{\prime}],$ (23)
@@
III Results and constraints
III.1 Leading Order Semi-Analytical Result
@@
To derive a simple analytic estimate, we focus instead on the leading $nlm=011$ state. In leading order in $\mu R_{\rm{LC}}$ we find
$\Gamma^{011}_{\text{SR}}$ $=Cg_{\rm{a}\gamma\gamma}^{2}B_{0}^{2}(\mu R)^{6}\mathcal{N}_{\phi}^{2}\alpha_{01}^{5}\frac{\sigma}{\mu}(\mu R_{\rm{LC}})^{5}(\Omega-\frac{170}{141}\mu),$ (25)
where $C=\frac{4.7\times 10^{-3}\pi^{2}}{8\Gamma(\frac{5}{2})^{2}}\approx 3.25\times 10^{-3}$, $\omega\approx\mu$ has been used and $R_{\rm{LC}}\equiv\Omega^{-1}$ is the light cylinder radius of the star where rigid rotation with the star is no longer possible. In the radial integral we model the conductivity as in [43], i.e. assume that $\sigma=$const. for $R\leq r\leq R_{\rm{LC}}$, $\sigma=\infty$ for $r<R$ and $\sigma=0$ for $r>R_{\rm LC}$.
In Fig. 1 we compare the leading-order semi-analytical rate of Eq. (25) and the numerical result, obtained by evaluating Eq. (18) using the matrix elements given in Eq. (23) and in Appendix A. We find that they agree remarkably well. For higher axion masses the relation slowly fails as the leading order $\alpha\propto\mu$ expansion of the radial integrals ceases to be accurate enough [^1] The gravitational coupling $\alpha$ is also proportional to the mass of the star. We have considered the most massive MSP known, PSR J0952-0607, for which $\alpha$ is maximal, to highlight that the semi-analytical formula always works. We note that the superradiance rate scales as $(\mu R)^{6}\alpha_{nl}^{2l+3}$. This has an extra factor of $\alpha_{nl}$ in relation to the result discussed in [43].
@@
$\rm{Im}(\Gamma_{31})$ $=-\frac{6}{50}\kappa^{2}\gamma\frac{\Omega}{\mu}\alpha_{01}^{2},\quad\rm{Im}(\Gamma_{32})=-\frac{24}{250}\kappa^{2}\gamma\frac{\Omega}{\mu}\alpha_{01}^{2},$
$\rm{Im}(\Gamma_{33})$ $=\frac{8}{50}\kappa^{2}\gamma(1-\frac{\Omega}{\mu})\alpha_{01}^{2},$ (32)
@@ III.2 Constraints
Figure 2: Bounds from the two largest spin MSPs known, as well as from a hypothetical 750Hz pulsar with a stiff equation of state. The solid black line represents the envelope of currently leading constraints [10].
The strongest constraints come from neutron stars which have long lifetimes and which have large masses and radii, as well as large magnetic fields. Of immediate interest are the fastest ($716$ Hz) millisecond pulsar PSR J1748-2446ad considered in Ref. [43] as well as the more recently observed heaviest ($2.35\ M_{\odot}$) and second fastest spinning ($707$ Hz) PSR J0952-0607 [53]. The mass and radius of PSR J1748-2446ad are not well known. As MSPs typically form by accretion, we conservatively take them to be $M=1.4\ M_{\odot}$ and $R=10$ km. For PSR J0952-0607 the mass and radius are known better [53], and we take them to be $M=2.35M_{\odot}$ and $R=12.5$ km. We assume the surface magnetic field to be $B_{0}=10^{8}$ G for both, and we take the conductivity in both cases to be $\sigma=10^{-12}$ eV, which is approximately in the middle of the observed range $0.01\Omega\lesssim\sigma\lesssim 100\Omega$ [49] for pulsars. Finally, we take a hypothetical 750 Hz MSP with mass and radius again being given $M=2.35M_{\odot}$ and $R=12.5$ km, but with slightly stronger magnetic field $5\times 10^{8}$ G. We take the Shklovskii-corrected period derivatives $\dot{P}$ needed to compute the characteristic lifetimes from the ATNF pulsar catalogue [52]. For PSR J1748-2446ad, PSR J0952-0607 and the hypothetical pulsar we use $\dot{P}=7.05\times 10^{-22}$, $\dot{P}=4.77\times 10^{-21}$ and $\dot{P}=10^{-22}$ respectively. The constraints from each are shown in Fig. 2 together with the current leading constraints from Ref. [10].
@@
IV Discussion and conclusions
@@
In this paper we have computed perturbatively the superradiance rate for a finite-conductivity instability in the plasma using a dipole magnetic field configuration. We have applied this to constrain the axion-photon coupling for axion mass $\sim 10^{-12}$ eV, using the fastest known MSP PSR J1748-2446ad and the supermassive PSR J0952–0607. Additionally we have given a projection for a hypothetical 750 Hz pulsar with a stiff equation of state. We see that known millisecond pulsars are unlikely to provide competitive bounds on axions. The hypothetical example shows, however, that if new fast MSPs are found with future radio telescopes, the situation could change.
@@
Note added
While this paper was in preparation, the effect of photon plasma mass on the instability was analyzed in Ref. [58]. They find that the instability is additionally suppressed by the plasma mass which is generally larger than the axion mass, $\omega_{p}>\mu$, weakening the prospects of constraining axions by superradiance in neutron star magnetospheres.
@@ Appendix B List of auxiliary functions
$f_{2}^{32}$ $={\sin^{2}{\theta}}{\cos{\phi}}e^{-i\phi}\qquad$ $f_{3}^{32}$ $=-{\sin{\theta}}{\cos{\theta}}{\sin{\phi}}$
@@ Appendix C Expansion of radial integrals
$I[l-a,l,l^{\prime}]\simeq$ $\pi{n+2l+1\choose n}2^{-l^{\prime}-1}\alpha^{l+1-a}\Bigg{(}\frac{\alpha\Gamma(l+\tilde{l}-a+2)}{\Gamma(\frac{1}{2}(-l+l^{\prime}+a+1))\Gamma(\frac{1}{2}(l+l^{\prime}-a+3))}$

### v2

@@ I Introduction
In a plasma background the photon obtains an effective mass $\omega_{p}$. If this mass is large enough, the superradiant scattering can be effectively blocked [49]. We discuss how $\omega_{p}$ is affected by the modeling of the magnetosphere. In this regard our main conclusion is that quantitative features of the effect of $\omega_{p}$ are very sensitive to the details of the dynamics and the superradiant mechanism we consider in this paper remains a viable possibility.
The paper is organized as follows: In Sec. II we briefly introduce the calculational framework we use and explicit definitions and formulas needed for the evaluation of the final results. In Sec. III we present our main results and apply them to constrain axion parameter space using observations of MSPs and in Sec. IV we discuss the conditions under which photon plasma mass can efficiently suppress superradiant scattering in the magnetosphere. In Sec. V we summarize and discuss further prospects of this work. Detailed intermediate results are provided in Appendices.
@@ II.2 Matrix elements for a dipole magnetic field
In Ref. [43] superradiance was demonstrated in the simple case of a magnetic field in the $\hat{\bf{z}}$-direction. This turns out to approximately capture the essential features of the system in the $l=|m|=1$ case, as the axion cloud is localized close to the equator, where the physical magnetic field is approximately in the $\hat{\bf{z}}$-direction. However, to extract more precise phenomenological predictions one must model the magnetic field more carefully. We therefore assume the more realistic dipole magnetic field
$\mathbf{B}(r,\theta,\phi)=B_{0}\big{(}\frac{R}{r}\big{)}^{3}(2\cos\theta\hat{\bm{r}}+\sin\theta\hat{\bm{\theta}}),$ (21)
@@
$\left\langle\phi_{lmn}\left|V_{a\gamma\gamma}\right|A_{l^{\prime}m^{\prime}}^{(i)}\right\rangle$ $=(-1)^{-m}g_{\rm{a}\gamma\gamma}B_{0}R^{3}\mathcal{N}_{\phi}(\mu\alpha_{nl}\omega)^{\frac{3}{2}}\mathcal{I}^{(i)}_{\Omega}I[l-2,l,l^{\prime}],$ (23)
@@
III Results
III.1 Leading Order Semi-Analytical Formula
@@
To derive a simple analytic estimate, we focus instead on the leading $nlm=011$ state. In leading order in $\mu R_{\rm{LC}}$ we find the superradiance rate to be
$\Gamma^{011}_{\text{SR}}$ $=\frac{\pi}{225}g_{\rm{a}\gamma\gamma}^{2}B_{0}^{2}(\mu R)^{6}\mathcal{N}_{\phi}^{2}\alpha_{01}^{5}\frac{\sigma}{\mu}(\mu R_{\rm{LC}})^{5}(\Omega-\frac{17}{15}\mu),$ (25)
where $\omega\approx\mu$ has been used and $R_{\rm{LC}}\equiv\Omega^{-1}$ is the light cylinder radius of the star where magnetic field lines can no longer rigidly rotate with the star. In the radial integral we model the conductivity as in [43], i.e. assume that $\sigma=$const. for $R\leq r\leq R_{\rm{LC}}$, $\sigma=\infty$ for $r<R$ and $\sigma=0$ for $r>R_{\rm LC}$.
In Fig. 1 we compare the leading-order semi-analytical rate of Eq. (25) and the numerical result, obtained by evaluating Eq. (18) using the matrix elements given in Eq. (23) and in Appendix A. We find that they agree remarkably well. For higher axion masses the relation slowly fails as the leading order $\alpha\propto\mu$ expansion of the radial integrals ceases to be accurate enough [^1] The gravitational coupling $\alpha$ is also proportional to the mass of the star. We have considered the most massive MSP known, PSR J0952-0607, for which $\alpha$ is maximal, to highlight that the semi-analytical formula is always valid.. We note that the superradiance rate scales as $(\mu R)^{6}\alpha_{nl}^{2l+3}$. This has an extra factor of $\alpha_{nl}$ in relation to the result discussed in [43].
@@
$\rm{Im}(\Gamma_{31})$ $=\rm{Im}(\Gamma_{32})=-\frac{6}{50}\kappa^{2}\gamma\frac{\Omega}{\mu}\alpha_{01}^{2},\quad\rm{Im}(\Gamma_{33})=\frac{8}{50}\kappa^{2}\gamma(1-\frac{\Omega}{\mu})\alpha_{01}^{2},$ (32)
@@ III.2 Constraints
Figure 2: Projected bounds from the two largest spin MSPs known, as well as from a hypothetical 750 Hz pulsar with a stiff equation of state and assuming negligible plasma effects. The solid black line represents the envelope of currently leading constraints [10].
The strongest constraints come from neutron stars which have long lifetimes and which have large masses and radii, as well as large magnetic fields. Of immediate interest are the fastest ($716$ Hz) millisecond pulsar PSR J1748-2446ad considered in Ref. [43] as well as the more recently observed heaviest ($2.35\ M_{\odot}$) and second fastest spinning ($707$ Hz) PSR J0952-0607 [54]. The mass and radius of PSR J1748-2446ad are not well known. As MSPs typically form by accretion, we conservatively take them to be $M=1.4\ M_{\odot}$ and $R=10$ km. For PSR J0952-0607 the mass and radius are known better [54], and we take them to be $M=2.35M_{\odot}$ and $R=12.5$ km. We assume the surface magnetic field to be $B_{0}=10^{8}$ G for both, and we take the conductivity in both cases to be $\sigma=10^{-12}$ eV, which is approximately in the middle of the observed range $0.01\Omega\lesssim\sigma\lesssim 100\Omega$  [50] for pulsars. Finally, we take a hypothetical 750 Hz MSP with mass and radius again being given $M=2.35M_{\odot}$ and $R=12.5$ km, but with slightly stronger magnetic field $5\times 10^{8}$ G. We take the Shklovskii-corrected period derivatives $\dot{P}$ needed to compute the characteristic lifetimes from the ATNF pulsar catalogue [53]. For PSR J1748-2446ad, PSR J0952-0607 and the hypothetical pulsar we use $\dot{P}=7.05\times 10^{-22}$, $\dot{P}=4.77\times 10^{-21}$ and $\dot{P}=10^{-22}$ respectively. The projected bounds from each are shown in Fig. 2 together with the current leading constraints from Ref. [10]. The bounds are valid under the assumption that there is no strong suppression due to the photon plasma mass. The conditions under which this can be expected to hold are discussed in detail in Sec. IV.
@@
IV Photon plasma mass and magnetosphere model
In a plasma background the photon obtains an effective mass $\omega_{p}$ which can be larger than the mass of an axion capable of stellar superradiance. In such case the axion-photon conversion probability $p_{a\rightarrow\gamma}=|\langle A|V_{a\gamma}|\phi\rangle|^{2}$ is suppressed by an additional factor of $\sim(\frac{\omega}{\omega_{p}})^{4}$. As argued in Ref. [49], this plasma mass can effectively block a superradiantly scattered photon from depositing its energy back into the axion sector. One might also worry that the photon plasma mass, related to the imaginary part of the conductivity, might affect the superradiance mechanism itself. In this section we will briefly address the possible suppressive effect of the photon plasma mass and argue that the imaginary part of the conductivity does not qualitatively change the superradiance picture. We also justify the use of a global dipole magnetic field in this work.
The photon plasma mass, assuming a standard Goldreich-Julian (GJ) model for the magnetosphere with $n_{e}=|\frac{\Omega\cdot B}{2\pi e}|$ and a dipole magnetic field (21), is given by
$\omega_{p}$ $=\sqrt{\frac{n_{\mathrm{e}}e^{2}}{m_{\mathrm{e}}}}=\sqrt{\frac{e\Omega}{2\pi}\frac{B_{0}}{m_{e}}\frac{R^{3}}{r^{3}}(3\cos^{2}{\theta}-1)},$ (36)
which for a typical millisecond pulsar is around $\sim 10^{-8}\sqrt{(3\cos^{2}{\theta}-1)}\mathrm{\,eV}$, reaching as low as $\sim 10^{-9}\mathrm{\,eV}$ at $\theta=53^{\circ}$ declination. Thus naively one would indeed expect superradiance to be strongly damped. However, we will now argue that already within the GJ model it is possible to significantly alleviate this suppression. Additionally, the GJ model may not always be a suitable to model millisecond pulsar magnetospheres to begin with, and that in other magnetospheric models the expected plasma suppression may even be negligible.
Assuming a power law scaling $B=B_{\ast}(\frac{R_{\ast}}{r})^{\gamma}$, a typical radial integral in an axion-photon matrix element, see Eq. (22), scales as
[…]
where the subscript $\ast$ is to be understood as the corresponding quantity at the surface of the star (for $r<R_{\rm{LC}}$) or at the light cylinder (for $r>R_{\rm{LC}}$). As the $\gamma$-scaling is generally different in the two regimes, one should evaluate the integral in two pieces with the above in mind. For the dipole field in Eq. (21), we have $\gamma=3$.
Numerically evaluating the integral in Eq. (37) we find that for the parameters relevant for superradiance, the radial integral contribution outside the light cylinder is generally dominant by factors of $\sim 10$ already for a dipole magnetic field ansatz regardless of plasma. Additionally, in the GJ model, the magnetic field can no longer corotate with the star outside the light cylinder, forming instead a pulsar wind which scales as $B\sim 1/r$ (i.e. $\gamma=1$) at large distances [56]. This further enhances the radial integral outside.
Inside the pulsar wind the electron density is typically assumed to fall as $n_{e}\sim 1/r^{2}$, under the assumption of particle number conservation, spherical symmetry, no collimation of the wind and no acceleration of particles. However, particles can be further accelerated in the pulsar wind to Lorentz factors up to $\sim 10^{5}$ or higher [57]. This acceleration occurs to distance scales much larger than those relevant for superradiance, with the Lorentz factor growing approximately linearly with distance [58]. As a consequence, the plasma density outside can fall off much faster. These considerations leave plausible the scenario of pulsar wind superradiance.
In the GJ model it is assumed that the whole magnetosphere is filled with plasma. However, for pulsars of periods shorter than a second and magnetic field strengths below $B<10^{9}$ G, pair-creation is not efficient enough to fill the whole magnetosphere [59]. Instead, there can be large regions of low charge density inside the pulsar magnetosphere. This occurs for instance in the electrosphere solution where charges are fully separated to domes of electrons around the polar caps and an equatorial torus of positrons and protons [59]. This model is more realistic than the naive force-free magnetosphere for millisecond pulsar magnetospheres, and is observed to arise naturally in simulations when pair production rate is low [60]. The distributions of all charged particles are shown concisely in Fig. 1 of Ref. [60]. We note here that around the equator where the axion cloud is localized, the number densities of electrons, positrons and protons are given approximately by $n_{\mathrm{e^{-}}}\approx n_{\mathrm{e^{+}}}\approx 0$ and $n_{\mathrm{p}}\sim 10^{-2}n_{GJ}$. Due to their large inertia, the protons do not contribute significantly to the photon plasma mass. Thus, we expect that neglecting the plasma mass is an excellent approximation in this scenario. Additionally, as again seen in Fig. 1 of Ref. [60], in the electrosphere solution, away from the poles, the magnetic field lines close even beyond the light cylinder and behave as a dipole field. No causality requirement is broken by this, as there is no plasma moving along the field lines. This supports the global use of the dipole magnetic field in this work. In reality the nature of the magnetosphere is presumably something between the plasma-filled and charge-separated extremes. Indeed, in a perfectly charge-separated magnetosphere there would be no spindown, describing a completely dead pulsar [61], and as such some plasma leakage is required.
Finally we briefly consider the impact of the imaginary part of the conductivity, connected to the photon plasma mass by $\sigma_{I}=\omega_{p}^{2}/\omega$, on the superradiance mechanism. We have verified that numerically the eigenfrequency correction of Eq. (18) is proportional to $i\sigma_{R}-\sigma_{I}$, so that the superradiance rate depends only on $\sigma_{R}$. This conclusion holds in the perturbative regime $\sigma_{I},\sigma_{R}\ll\mu$, and for larger $\sigma_{I}$ one would expect some damping [42, 43]. As noted in Ref. [49], in the strongly-collisional regime $\omega\ll\nu$, with $\nu$ the collision rate in the Drude model, $\sigma_{I}\ll\sigma_{R}$ generally holds. This is true for instance in accreting pulsars in low-mass X-ray binaries.
V Discussion and conclusions
@@
In this paper we have computed perturbatively the superradiance rate for a finite-conductivity instability in the plasma using a dipole magnetic field configuration. We have applied this to constrain the axion-photon coupling for axion mass $\sim 10^{-12}$ eV, using the fastest known MSP PSR J1748-2446ad and the supermassive PSR J0952–0607. Additionally we have given a projection for a hypothetical 750 Hz pulsar with a stiff equation of state. We find that known millisecond pulsars are unlikely to provide competitive bounds on axions. The hypothetical example shows, however, that if new fast MSPs are found with future radio telescopes, the situation could change. It has been recently noted in Ref. [49], that the superradiance rate could be additionally suppressed by the photon plasma mass. We have discussed how the magnitude of this effect is strongly dependent on the detailed structure of the pulsar magnetosphere: for force-free case the suppression due to the plasma mass is significant, while for the charge-separated magnetosphere we find that the effect is small. The reality lies in between and remains to be determined more precisely.
@@ Appendix B List of auxiliary functions
$f_{2}^{32}$ $={\sin^{2}{\theta}}{\cos{\phi}}e^{-i\phi}\qquad$ $f_{3}^{32}$ $=-{\sin{\theta}}{\cos{\theta}}{\cos{\phi}}$
@@ Appendix C Expansion of radial integrals
$I[l-a,l,l^{\prime}]\simeq$ $\pi{n+2l+1\choose n}2^{-l^{\prime}-1}\alpha^{l+1-a}\Bigg{(}\frac{\alpha\Gamma(l+l^{\prime}-a+2)}{\Gamma(\frac{1}{2}(-l+l^{\prime}+a+1))\Gamma(\frac{1}{2}(l+l^{\prime}-a+3))}$